For the following set of cash flows,

  

Year Cash Flow
0 –$9,000           
1 6,100           
2 6,500           
3 5,000           

  

a. What is the NPV at a discount rate of 0 percent?

  

b. What is the NPV at a discount rate of 9 percent?

  

c. What is the NPV at a discount rate of 25 percent?
d. What is the NPV at a discount rate of 30 percent?
Answer #1

a.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=6100+6500+5000

=17600

NPV=Present value of inflows-Present value of outflows

=17600-$9000

=8600

b.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=6100/1.09+6500/1.09^2+5000/1.09^3

=14928.17

NPV=Present value of inflows-Present value of outflows

=14928.17-$9000

=5928.17(Approx).

c.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=6100/1.25+6500/1.25^2+5000/1.25^3

=11600

NPV=Present value of inflows-Present value of outflows

=11600-$9000

=2600

d.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=6100/1.3+6500/1.3^2+5000/1.3^3

=10814.29

NPV=Present value of inflows-Present value of outflows

=10814.29-$9000

=$1814.29(Approx).